OpenStudy (anonymous):

may you solve it? limx→02x2/1−cosx

myininaya (myininaya):

Ok ok this will try to write the bottom with one term by multiplying by bottom's conjugate We still need to use those two limits (or at least of the two) for this one also

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{2x^2}{1-\cos(x)}\] So we multiply (1+cos(x)) on top and bottom

OpenStudy (anonymous):

ok freind

OpenStudy (anonymous):

well...

OpenStudy (anonymous):

not a pal of l'hopital i guess

myininaya (myininaya):

lol thats a good guy to use here but i like the manipulations ;)

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{2x^2 \cdot (1+\cos(x))}{(1-\cos(x))(1+\cos(x))}\]

myininaya (myininaya):

that is the first step I was asking you to do so you can write the bottom as one term

myininaya (myininaya):

but if you prefer sat's way I will let him by himself with you

OpenStudy (anonymous):

fine. while you struggle i will write \[\lim_{x\to 0}\frac{2x^2}{1-\cos(x)}=\lim_{x\to 0}\frac{4x}{\sin(x)}=\lim_{x\to 0}\frac{4}{\cos(x)}\]

OpenStudy (anonymous):

and then wait for the algebra

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{2x^2 (1+\cos(x))}{\sin^2(x)}=2 \lim_{x \rightarrow 0}\frac{x^2}{\sin^2(x)}(1+\cos(x))\]

myininaya (myininaya):

What does x/sin(x)->? as x->0

OpenStudy (anonymous):

x/2

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{x}{\sin(x)}=1 \text{ by squeeze thm}\]

OpenStudy (anonymous):

actually this is pretty snappy isn't it?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}(\frac{x}{\sin(x)})^2=1^2=1\]

myininaya (myininaya):

For next part (1+cos(x)) This part is continuous at x=0 and se therefore you can just plug in x=0

myininaya (myininaya):

so then you have 2*1*( ?)

OpenStudy (anonymous):

satellite73 You must have sense

myininaya (myininaya):

What does that mean?

OpenStudy (anonymous):

I wasn't with you

OpenStudy (anonymous):

I was with her

OpenStudy (anonymous):

well... You say..